Bill traveled to town hall and back. It took three hours longer to go there than it did to come back. The average speed on the trip was 30 mph. The average speed on the way back was 75 mph. How many hours did the trip take?

Answer by lwsshak3(11628)

Bill traveled to town hall and back. It took three hours longer to go there than it did to come back. The average speed on the trip was 30 mph. The average speed on the way back was 75 mph. How many hours did the trip take?
***
let x=travel time of return trip
x+3=travel time of trip to town
distance=travel time*speed
..
30(x+3)=75x
30x+90=75x
45x=90
x=2
x+3=5
How many hours did the trip take? 5+2=7